OpenStudy (anonymous):

A yardstick measures 1/4 by 3 by 36 inches. How many yard sticks will fit in a box 3 inches wide and 36 inches high, if the girth of the box is 30 inches?

OpenStudy (anonymous):

Girth = 30 inches Girth =3 + 3 + x + x 30 = 6 +2x 2x=30-6=24 x=12 yardsticks are 1/4 in thick 12/.25=48 48 yardsticks could be packed in that box. I'm I right ? If you can check my work...

OpenStudy (anonymous):

27 yard sticks, not sure though If the box is 3 in wide and its girth is 30 in then its depth must be 12in. 1/4in thick means you'll fit 12 into 3in 3in wide means you'll get 4 in 12in so number of rulers = 12x4=48 A yardstick measures 1/4 by 3 by 36 inches. How many yard sticks will fit in a box 3 inches wide and 36 inches high, if the girth of the box is 30 inches?